# Navigating Cookie Consent Violations Across the Globe

### v1

@@ 6.4 Cookie Consent Violations Across Regions
Overall, the EU and UK have fewer undeclared cookies (87,251 and 54,633), websites with undeclared cookies (90.19% and 90.04%), and undeclared cookie consent violations as a proportion of all cookies. Conversely, the number of ignored cookie rejection violations when normalized to the total cookie count is higher in these regions (Figure 5). Such a phenomenon can be explained by the increase in proper cookie declarations. For a cookie to be classified as an ignored cookie rejection by our system, it needs to be declared by the cookie library. Thus, the increase in declared cookies results in more cookies being continued to be placed even after the user rejects all cookies. In contrast, the US had the most websites with undeclared cookies with 95.93% of websites containing at least one undeclared cookie violation. AU had the third-fewest cookies loaded in our measurements. The region also had the fewest websites with ignored cookie rejection violations with only 80.74% containing these violations. CA had the most websites with at least one ignored cookie rejection, with 88.41%. Across all regions, wrong cookie category violations were found on very few websites, between only 2.50% and 3.69% of websites had one violation of this type. Despite Canada’s PIPEDA having similar guidelines as the EU/UK regarding tracking cookie consent, this region contained more cookie consent violations. nhl.com contains cookies from googleadservices.com, an example of such an undeclared cookie violation (for tracking cookies in Canada). We found from our measurements that the same websites had many more cookies and cookie consent violations when accessed in different regions. Websites in non-EU/UK regions had roughly 8 additional cookies (Figure 8) and 7 additional cookie consent violations (Figure 9). As an example, nvidia.com had 16.25–24.00 cookies in AU, SG, EU, UK, CAN whereas it had 28.86–36.78 cookies in CA, ZA, US.
@@
Third-party cookies were most prevalent in the US across websites by a large margin (29.03 cookies per site). Regions with websites containing the fewest set of third-party cookies with only an average of 8.88 (EU) and 9.16 (UK) cookies per site. Other regions like CAN (22.87), AU (21.3), CA (20.26), SG (19.86), and ZA (19.52) had fewer third-party cookies loaded per website, but still significantly more than websites accessed from the EU and UK. These discrepancies may be attributed to a combination of advertisers’ interest in the regions’ users and region-specific privacy laws. CMPs provide website developers with templates and location rulesets to customize cookie banners and cookie placement depending on the region. Advertisers may have a higher interest in CA users for these English-speaking sites than regions in Asia or Africa (so cookies are more prevalent). Combined with CA’s stricter requirements on opt-outs, personal information, and cookie notices, website developers may declare more cookies for the CA region, but place fewer on these users’ browsers. North American regions have the most undeclared cookies, with CA containing the fewest. Whereas AU, SG, and ZA contain more undeclared cookies than the EU and UK. An illustrative example of this is sportinglife.com which uses 3rd-party cookies from Google and DoubleClick. In CA, the website loads these undeclared cookies (e.g., NID, IDE) from both regions, whereas in the EU, the website does not load these cookies. Generally, the disparities observed in cookie consent violations (Figure 5) are exacerbated in 3rd-party cookie consent violations (Figure 6).
@@ 7 Root Cause Analysis
We analyzed Onetrust and Cookiebot’s documentation and demos to gain insight into whether CMPs or website developers hold responsibility for cookie consent violations. Although CMPs may act as data controllers or processors [santos2021consent], both parties share responsibility. CMPs need to provide stronger guardrails, easier integration methods, support for more legal frameworks, and clearer documentation. Site owners need to ensure no undeclared cookies remain, properly implement cookie/script-blocking, and keep their CMP integrations updated. We find that most undeclared cookies result from either website developers not categorizing cookies or CMP crawlers missing some cookies, whereas ignored cookie rejections stem from developers not integrating CMP scripts correctly.
@@ 7.1 CMP Responsibility
Template Gaps: Most CMPs ship with templates only for major legal frameworks like GDPR and CCPA, leaving other jurisdictions unsupported. As a result, developers have no pre-configurations or guidelines for regions like Canada, Singapore, or South Africa, leading to inconsistent or noncompliant cookie banners (Figure 17). Note that these templates are useful for developers (unfamiliar with privacy laws) in determining cookie banner behaviors (e.g., whether they are shown, reconsent expiration time, etc. Figure 18).
@@
Geolocation Rulesets: CMPs offer geolocation rulesets that dynamically adjust banner text, button labels, consent models (opt-in vs. implied), and even hide “Reject All” to maximize opt-in rates per region (Figure 19). CMPs support geolocation rulesets, which can allow for customizing cookie banner stylization and functionality depending on the region of the user’s IP address. For example, the CCPA template for Onetrust supports translating behaviors like clicking or moving to the next page as the consent choice "Accept All Cookies" (Figure 18). At other locations without legal template support, website developers may set geolocation rulesets to not display cookie banners, or even allow scrolling to be counted as accepting all cookies. These rulesets account for the disparity in cookie placement and consent functionality between regions.
@@ 7.2 Developer Responsibility
Incorrect Script Embedding: In order for CMPs to block 3rd-party scripts that load cookies, website developers need to update any <script> tags with additional attributes (e.g., data-usercentrics="Google Maps" or <script class="optanon-category-C0002">). Additionally, CMP scripts connected to the Onetrust and Cookiebot services must be placed in the <head> before any other scripts [ashlea_cartee_say_2019, cybot_automatic_2019, cookiepro_onetrust_2021]. If these instructions are not followed perfectly, scripts will continue to load and cookies will be placed. Both OneTrust and Cookiebot’s documentation states that if cookies or tracking scripts are loaded before consent withdrawal, then the developer or user needs to deactivate the service or perform a page reload [onetrust2024docs, cookiebot2023docs]. A manual inspection revealed that websites left scripts uncontrolled by the CMPs. For example, on scientificamerican.com, we rejected Performance cookies, but found that both first- and third-party Analytics cookies such as Google Analytics (_ga) were still being loaded (Table 9).
Manual Cookie Categorization: When CMP auto-scanners miss new or previously uncategorized cookies, developers need to manually assign them to categories. In practice, these manual steps could be easily missed or skipped, so these “unknown” cookies remain undeclared and active. These missed cookies likely make up the bulk of the undeclared cookie violations (Figure 16).
Neglecting Updates: Websites frequently change [adar2009web], and web developers may forget to have CMPs re-crawl their sites and update their scripts, especially on new subdomains. The Cookiebot scanner is also priced to scale with the number of subpages to crawl, and thus may provide incomplete scans. These may also contribute to undeclared cookies.
@@ 10 Compliance with Open Science Policy
To uphold principles of transparency and reproducibility, the source code of our crawling and measurement system, alongside the analysis scripts and cookie data are available at https://github.com/byron123t/cookie-consent.
@@ 11 Acknowledgements
[…]
% Cookies CA EU UK AU US SG CAN ZA
Trackers 63.87% 64.10% 64.14% 63.36% 64.73% 62.89% 64.11% 64.40%
Location 7.34% 7.44% 7.38% 7.93% 8.18% 7.99% 8.19% 7.46%
IP Address 3.28% 5.74% 5.58% 3.51% 2.81% 3.54% 3.01% 3.81%
Language 0.54% 1.00% 0.92% 0.52% 0.45% 0.55% 0.51% 0.57%
Unlikely P.I. 24.98% 21.98% 21.30% 23.84% 25.03% 25.50% 24.09% 24.76%
Table 7: Violations and cookies containing personal information in the form of trackers, ip, or location data.
Cookie Name # Websites (%)
_ga 583 (48.4)
_gid 429 (35.6)
_fbp 318 (26.4)
IDE 296 (24.6)
_uetsid 218 (18.1)
Table 8: Top-5 Rejected Cookies Used.
Tracker # Websites (%)
doubleclick.net 328 (27.2)
linkedin.com 198 (16.4)
youtube.com 198 (16.4)
bing.com 88 (7.3)
twitter.com 69 (5.7)
Table 9: Top trackers of rejected-usage cookies.
Figure 8: Relative website-level differences in total cookies. (Compared to EU baseline)
Figure 9: Relative website-level differences in cookie violations. (Compared to EU baseline)
Figure 10: Reject all button presence in cookie banners.
Figure 11: Pairwise website-website cookie count differences.
Figure 12: Pairwise cookie consent lifetime differences.
Figure 13: Pairwise CMP behavior differences.
Figure 14: Distribution of labels of cookie banner buttons.
@@ G Feature (Dimension) HTML attributes $D_{G}$
$G_{1}$ n-grams and keywords (3) aria-label, class, id, text 12
$G_{2}$ tokens $>n_{t}$ or not (1) aria-label, text 2
@@ A.1 Button Extractor Data Collection
We randomly selected 8k websites from the top 10k global websites for developing features and training the button extractor. The remaining 2k websites were set aside to evaluate the model performance. We use the Tranco list [LePochat_tranco_2019] generated in July 2021 (ID: 9QK2) and accessed websites from an IP address in the UK to maximize encounters with cookie banners and preference buttons. We manually visited the home pages of 1,000 randomly selected websites from the 8k websites to identify cookie banner buttons. Two annotators independently annotated the home pages of the websites. For each website, a snapshot of the home page HTML and the CSS selectors were recorded. Only the websites with English home pages were annotated, excluding non-English pages and duplicates. We obtained a training set of 298 web pages containing 436 cookie banner buttons out of 71,020 all links/buttons. Many websites only show a cookie banner without any choice or only a binary accept/reject option.
@@ A.2 Button Extractor Feature Selection
We derive 3 classification feature groups based on the HTML attributes: aria-label, class, id, and inner text. The attribute aria-label, an accessibility feature of the web for marking buttons with labels for users with disabilities [cookiebot_functions_2021], is especially useful in cases where buttons are displayed as non-textual icons. The features are shown in Table 10 and the feature vector has 17 dimensions in total. Feature group $G_{1}$ is the number of occurrences of selected unigrams, bigrams and high-frequency keywords in the button labels. We also separate the most frequently-used bigrams into a set of high-frequency keywords (Table 11 lists some examples). Feature group $G_{2}$ indicates whether or not the number of tokens of a button label is greater than a threshold $n_{t}=9$ (empircally selected using training set). This feature allows our classifier to avoid long paragraphs that may contain keywords. Finally, feature group $G_{3}$ indicates whether a cookie consent library API is being used.
@@ A.3 Button Extractor Performance Metric
We report the recall@k score, a metric used to evaluate information retrieval systems [polisis_usenix_2018, thung2013automated, sun2011towards], which represents the portion of websites containing cookie banner buttons detected from the top k classification results. A website is successfully detected if one of its cookie banner buttons is among the top k buttons with the highest classification probabilities.
@@ Appendix B Consent Cookie Decoding
To extract cookie consent preferences, we decode consent cookies basing on the documentation and analyzing their key-value pairs. OneTrust’s consent cookie is called OptanonConsent [onetrust_llc_onetrust_2021] which stores the consent preference of each cookie category. For example, groups=C1:1,C2:0 indicates that cookie category $C1$ is approved while $C2$ is rejected. Cookiebot’s consent cookie is called CookieConsent storing consents for 4 fixed cookie categories: Necessary, Preferences, Statistics, and Marketing [cybot_developer_2021]. Similar to ’Necessary’ cookies, ’Unclassified’ cookies are not automatically blocked and cannot be denied by users, so the consent preferences for these cookies are by default set to True [cybot_cookiebot_2021].

### v2

@@ 6.4 Cookie Consent Violations Across Regions
Overall, the EU and UK have fewer undeclared cookies (87,251 and 54,633), websites with undeclared cookies (90.19% and 90.04%), and undeclared cookie consent violations as a proportion of all cookies. Conversely, the number of ignored cookie rejection violations when normalized to the total cookie count is higher in these regions (Figure˜5). Such a phenomenon can be explained by the increase in proper cookie declarations. For a cookie to be classified as an ignored cookie rejection by our system, it needs to be declared by the cookie library. Thus, the increase in declared cookies results in more cookies being continued to be placed even after the user rejects all cookies. In contrast, the US had the most websites with undeclared cookies with 95.93% of websites containing at least one undeclared cookie violation. AU had the third-fewest cookies loaded in our measurements. The region also had the fewest websites with ignored cookie rejection violations with only 80.74% containing these violations. CA had the most websites with at least one ignored cookie rejection, with 88.41%. Across all regions, wrong cookie category violations were found on very few websites, between only 2.50% and 3.69% of websites had one violation of this type. Despite Canada’s PIPEDA having similar guidelines as the EU/UK regarding tracking cookie consent, this region contained more cookie consent violations. nhl.com contains cookies from googleadservices.com, an example of such an undeclared cookie violation (for tracking cookies in Canada). We found from our measurements that the same websites had many more cookies and cookie consent violations when accessed in different regions. Websites in non-EU/UK regions had roughly 8 additional cookies (Figure˜8) and 7 additional cookie consent violations (Figure˜9). As an example, nvidia.com had 16.25–24.00 cookies in AU, SG, EU, UK, CAN whereas it had 28.86–36.78 cookies in CA, ZA, US.
@@
Third-party cookies were most prevalent in the US across websites by a large margin (29.03 cookies per site). Regions with websites containing the fewest set of third-party cookies with only an average of 8.88 (EU) and 9.16 (UK) cookies per site. Other regions like CAN (22.87), AU (21.3), CA (20.26), SG (19.86), and ZA (19.52) had fewer third-party cookies loaded per website, but still significantly more than websites accessed from the EU and UK. These discrepancies may be attributed to a combination of advertisers’ interest in the regions’ users and region-specific privacy laws. CMPs provide website developers with templates and location rulesets to customize cookie banners and cookie placement depending on the region. Advertisers may have a higher interest in CA users for these English-speaking sites than regions in Asia or Africa (so cookies are more prevalent). Combined with CA’s stricter requirements on opt-outs, personal information, and cookie notices, website developers may declare more cookies for the CA region, but place fewer on these users’ browsers. North American regions have the most undeclared cookies, with CA containing the fewest. Whereas AU, SG, and ZA contain more undeclared cookies than the EU and UK. An illustrative example of this is sportinglife.com which uses 3rd-party cookies from Google and DoubleClick. In CA, the website loads these undeclared cookies (e.g., NID, IDE) from both regions, whereas in the EU, the website does not load these cookies. Generally, the disparities observed in cookie consent violations (Figure˜5) are exacerbated in 3rd-party cookie consent violations (Figure˜6).
@@ 7 Root Cause Analysis
We analyzed Onetrust and Cookiebot’s documentation and demos to gain insight into whether CMPs or website developers hold responsibility for cookie consent violations. Although CMPs may act as data controllers or processors [58], both parties share responsibility. CMPs need to provide stronger guardrails, easier integration methods, support for more legal frameworks, and clearer documentation. Site owners need to ensure no undeclared cookies remain, properly implement cookie/script-blocking, and keep their CMP integrations updated. We find that most undeclared cookies result from either website developers not categorizing cookies or CMP crawlers missing some cookies, whereas ignored cookie rejections stem from developers not integrating CMP scripts correctly.
@@ 7.1 CMP Responsibility
Template Gaps: Most CMPs ship with templates only for major legal frameworks like GDPR and CCPA, leaving other jurisdictions unsupported. As a result, developers have no pre-configurations or guidelines for regions like Canada, Singapore, or South Africa, leading to inconsistent or noncompliant cookie banners (Figure˜17). Note that these templates are useful for developers (unfamiliar with privacy laws) in determining cookie banner behaviors (e.g., whether they are shown, reconsent expiration time, etc. Figure˜18).
@@
Geolocation Rulesets: CMPs offer geolocation rulesets that dynamically adjust banner text, button labels, consent models (opt-in vs. implied), and even hide “Reject All” to maximize opt-in rates per region (Figure˜19). CMPs support geolocation rulesets, which can allow for customizing cookie banner stylization and functionality depending on the region of the user’s IP address. For example, the CCPA template for Onetrust supports translating behaviors like clicking or moving to the next page as the consent choice "Accept All Cookies" (Figure˜18). At other locations without legal template support, website developers may set geolocation rulesets to not display cookie banners, or even allow scrolling to be counted as accepting all cookies. These rulesets account for the disparity in cookie placement and consent functionality between regions.
@@ 7.2 Developer Responsibility
Incorrect Script Embedding: In order for CMPs to block 3rd-party scripts that load cookies, website developers need to update any <script> tags with additional attributes (e.g., data-usercentrics="Google Maps" or <script class="optanon-category-C0002">). Additionally, CMP scripts connected to the Onetrust and Cookiebot services must be placed in the <head> before any other scripts [5, 21, 20]. If these instructions are not followed perfectly, scripts will continue to load and cookies will be placed. Both OneTrust and Cookiebot’s documentation states that if cookies or tracking scripts are loaded before consent withdrawal, then the developer or user needs to deactivate the service or perform a page reload [3, 19]. A manual inspection revealed that websites left scripts uncontrolled by the CMPs. For example, on scientificamerican.com, we rejected Performance cookies, but found that both first- and third-party Analytics cookies such as Google Analytics (_ga) were still being loaded (Table 9).
Manual Cookie Categorization: When CMP auto-scanners miss new or previously uncategorized cookies, developers need to manually assign them to categories. In practice, these manual steps could be easily missed or skipped, so these “unknown” cookies remain undeclared and active. These missed cookies likely make up the bulk of the undeclared cookie violations (Figure˜16).
Neglecting Updates: Websites frequently change [2], and web developers may forget to have CMPs re-crawl their sites and update their scripts, especially on new subdomains. The Cookiebot scanner is also priced to scale with the number of subpages to crawl, and thus may provide incomplete scans. These may also contribute to undeclared cookies.
@@ 10 Compliance with Open Science Policy
To uphold principles of transparency and reproducibility, the source code of our crawling and measurement system, alongside the analysis scripts and cookie data are available at https://github.com/byron123t/cookie-consent or https://doi.org/10.5281/zenodo.15566975.
@@ G Feature (Dimension) HTML attributes $D_{G}$
$G_{1}$ # n-grams and keywords (3) aria-label, class, id, text 12
$G_{2}$ # tokens $n_{t}$ or not (1) aria-label, text 2
@@ A.1 Button Extractor Data Collection
We randomly selected 8k websites from the top 10k global websites for developing features and training the button extractor. The remaining 2k websites were set aside to evaluate the model performance. We use the Tranco list [42] generated in July 2021 (ID: 9QK2) and accessed websites from an IP address in the UK to maximize encounters with cookie banners and preference buttons. We manually visited the home pages of 1,000 randomly selected websites from the 8k websites to identify cookie banner buttons. Two annotators independently annotated the home pages of the websites. For each website, a snapshot of the home page HTML and the CSS selectors were recorded. Only the websites with English home pages were annotated, excluding non-English pages and duplicates. We obtained a training set of 298 web pages containing 436 cookie banner buttons out of 71,020 all links/buttons. Many websites only show a cookie banner without any choice or only a binary accept/reject option.
@@ A.2 Button Extractor Feature Selection
We derive 3 classification feature groups based on the HTML attributes: aria-label, class, id, and inner text. The attribute aria-label, an accessibility feature of the web for marking buttons with labels for users with disabilities [17], is especially useful in cases where buttons are displayed as non-textual icons. The features are shown in Table 10 and the feature vector has 17 dimensions in total. Feature group $G_{1}$ is the number of occurrences of selected unigrams, bigrams and high-frequency keywords in the button labels. We also separate the most frequently-used bigrams into a set of high-frequency keywords (Table 11 lists some examples). Feature group $G_{2}$ indicates whether or not the number of tokens of a button label is greater than a threshold $n_{t}=9$ (empircally selected using training set). This feature allows our classifier to avoid long paragraphs that may contain keywords. Finally, feature group $G_{3}$ indicates whether a cookie consent library API is being used.
@@ A.3 Button Extractor Performance Metric
We report the recall@k score, a metric used to evaluate information retrieval systems [34, 66, 61], which represents the portion of websites containing cookie banner buttons detected from the top k classification results. A website is successfully detected if one of its cookie banner buttons is among the top k buttons with the highest classification probabilities.
@@ Appendix B Consent Cookie Decoding
To extract cookie consent preferences, we decode consent cookies basing on the documentation and analyzing their key-value pairs. OneTrust’s consent cookie is called OptanonConsent [49] which stores the consent preference of each cookie category. For example, groups=C1:1,C2:0 indicates that cookie category $C1$ is approved while $C2$ is rejected. Cookiebot’s consent cookie is called CookieConsent storing consents for 4 fixed cookie categories: Necessary, Preferences, Statistics, and Marketing [23]. Similar to ’Necessary’ cookies, ’Unclassified’ cookies are not automatically blocked and cannot be denied by users, so the consent preferences for these cookies are by default set to True [22].